### v4

# A universal speed limit for spreading of coherence

Abstract
Discoveries of fundamental limits for the rates of physical processes, from the speed of light to the Lieb–Robinson bound for information propagation [1, 2], often lead to breakthroughs in the understanding of the underlying physics. Here we observe such a limit for a paradigmatic many-body phenomenon, the spreading of coherence during formation of a weakly interacting Bose–Einstein condensate [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. We study condensate formation in an isolated homogeneous atomic gas [19, 20] that is initially far from equilibrium, in an incoherent low-energy state, and condenses as it relaxes towards equilibrium. Tuning the inter-atomic interactions that drive condensation, we show that the spreading of coherence through the system is initially slower for weaker interactions, and faster for stronger ones, but always eventually reaches the same limit, where the square of the coherence length grows at a universal rate given by the ratio of Planck’s constant and the particle mass, or equivalently by the quantum of velocity circulation associated with a quantum vortex. These observations are robust to changes in the initial state, the gas density, and the system size. Our results provide benchmarks for theories of universality far from equilibrium [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], are relevant for quantum technologies that rely on large-scale coherence, and invite similar measurements in other systems.
Understanding the dynamics of far-from-equilibrium many-body systems, including the emergence of long-range order in such systems, is an outstanding problem in physics, relevant from subnuclear to cosmological lengthscales [21, 22, 23, 24, 26, 25, 27, 28, 29, 30, 31, 32, 33, 34]. Conceptually, far-from-equilibrium relaxation and emergence of coherence have long been linked to decaying turbulence [21, 23, 12, 27], which often features self-similar scaling dynamics. More recently, within the framework of nonthermal fixed points (NTFPs) [28], theorists have drawn analogies between such dynamics and the equilibrium properties of systems close to a continuous phase transition. Near the transition to an ordered state of matter, such as a superfluid or a ferromagnet, the system is scale invariant and its salient properties do not depend on the microscopic details [35]. Analogously, in the NTFP theory, far-from-equilibrium systems, including the early universe undergoing reheating [27], quark-gluon plasma in heavy-ion collisions [36], quantum magnets [37], and ultracold atomic gases [38, 39, 40, 41, 42, 43], generically display dynamic (spatiotemporal) scaling, with scaling exponents that could define far-from-equilibrium universality classes. Recently, far-from-equilibrium dynamic scaling was observed in several experiments with ultracold atoms, in both isolated (relaxing) [44, 45, 46, 47, 48, 49, 50] and continuously driven [51, 52] systems.
Here, we go beyond the elegant scaling properties of far-from-equilibrium relaxation [44, 45, 46, 47, 48, 49, 50], to address the crucial question of how long it takes to establish long-range order. We study this problem for the paradigmatic macroscopically coherent state, the weakly interacting Bose–Einstein condensate, which is also a textbook example of a superfluid.
We study condensate formation in an isolated homogeneous Bose gas [53], trapped in a cylindrical optical box [19, 20] such as sketched in Fig. 1a (Methods). The gas is prepared far from equilibrium and is initially incoherent, but has very low energy and condenses as it relaxes towards equilibrium under the influence of inter-atomic interactions, characterised by the $s$-wave scattering length $a$. As illustrated in Fig. 1a, in a homogeneous system the (global) condensate grows through coarsening [25], the local spreading of coherence. This coarsening is quantified by the growth of the coherence length $\ell$, over which the first-order correlation function $g_{1}(r)$ decays, and corresponds to narrowing of the momentum distribution $n_{k}(k)$ (where $k$ is the wavenumber), which is related to $g_{1}(r)$ by the Fourier transform.
Our experiments are performed with $^{39}\textrm{K}$ atoms, and we tune $a$ using a Feshbach resonance, exploring coarsening for $a=(50-400)\,a_{0}$, where $a_{0}$ is the Bohr radius. Our cylindrical box has radius $R=21(2)\,\upmu{\rm m}$, length $L=40(4)\upmu{\rm m}$, and volume $V=55(12)\times 10^{3}\,\upmu{\rm m}^{3}$. Our gas density $n=5.4(1.2)\,\upmu{\rm m}^{-3}$ corresponds (assuming ideal-gas thermodynamics) to the critical temperature for condensation $T_{\textrm{c}}=127(19)\,\textrm{nK}$. The kinetic energy per particle in our initial incoherent states is $\varepsilon=k_{\textrm{B}}\times 20(2)\,\textrm{nK}$ (where $k_{\textrm{B}}$ is the Boltzmann constant), corresponding to a large equilibrium condensed fraction $\eta=0.61(4)$ (see Extended Data Figs. 1 and 2). During coarsening the total particle number, $N\approx 3\times 10^{5}$, is essentially constant (see Methods), and the gas is always weakly interacting in the sense that $na^{3}<10^{-4}$. We measure $n_{k}$ by absorption imaging (along the $z$ direction) after time-of-flight expansion, performing the inverse-Abel transform on the line-of-sight integrated distributions; just for the images shown in Figs. 1b and  2a we instead image only slices of the cloud [54] corresponding to $k_{z}\approx 0$ (Methods). We normalise $n_{k}$ such that $\int n_{k}\,4\pi k^{2}\textrm{d}k=N$.
Figure 1: Universal coarsening of an isolated Bose gas. a, Real-space cartoon of coarsening. b, Momentum-space relaxation for different far-from-equilibrium initial states. Our initial states P1,2,3 (left column) have different momentum distributions $n_{k}$, but the same energy, so the gas always relaxes towards the same equilibrium state. For P1,2,3, the system takes different times, $t_{1,2,3}$, to evolve to the same $n_{k}$ shown in the middle column, but from this point onwards it always evolves in the same way. The $n_{k}$ distributions are averages of at least 2020 measurements. The red scale bar (top image) shows $1\,\upmu{\rm m}^{-1}$. c, Growth of the coherence length, $\ell$ (see text). Plotting $\ell^{2}(t)$ reveals three stages of relaxation: (i) the non-universal initial dynamics, (ii) the scaling regime where $\ell^{2}$ grows linearly (dashed lines), as expected for the scaling exponent $\beta=1/2$, and (iii) the breakdown of scaling at long times due to finite-size effects. The curves for P1,2,3 are parallel, with the initial-state effects captured by the different time offsets $t^{*}$ (intercepts of the dashed lines). d, Dynamic scaling. In the scaling regime, the full low-$k$ distributions for all three initial states (left panel) can be collapsed onto the same curve (right panel) according to Eqs. (2, 3) with $\beta=1/2$ and $t\rightarrow t_{\textrm{uni}}\equiv t-t^{*}$ (see also Extended Data Fig. 4); $t_{0}=60\,\textrm{ms}$ is an arbitrary reference time. All error bars show standard errors of the measurements.
We first show that, while the short-time relaxation dynamics inevitably depend on the details of the initial state, the long-time relaxation does not (Fig. 1b). For this purpose, we engineer three different far-from-equilibrium states, starting with a quasi-pure condensate and using a time-varying force to perturb the cloud (see Extended Data Fig. 1). Our initial states P1,2,3 have different $n_{k}$ (see left column), but the same $\varepsilon$. At time $t=0$ the gas is non-interacting and we then initiate relaxation by switching $a$ to $100\,a_{0}$. Starting from an initial state, $n_{k}(k)$ follows some trajectory (represented by the wiggly coloured lines) in the space of functions with the same $N$ and $\varepsilon$. The middle column shows that, still far from equilibrium, these trajectories converge to the same $n_{k}$. The time that the gas takes to evolve to this $n_{k}$ depends on the initial state ($t_{i=1,2,3}$ for Pi=1,2,3), but further evolution from this $n_{k}$ is the same for all initial states; note that the state trajectories for P1 and P2 merge before merging with the P3 one, but we just show an $n_{k}$ where all three have converged.
Figure 2: Spreading of coherence for different interaction strengths. a, Gas evolution for two different scattering lengths $a$, starting in the same initial state P1. For stronger interactions, the condensate (the peak at $k=0$) emerges sooner. b, However, plotting $\ell^{2}(t)$ reveals that the interaction strength affects only the non-universal initial dynamics (open symbols), while the linear growth of $\ell^{2}$ in the universal coarsening regime (solid symbols) is the same for both $a$ values.
The long-time relaxation of a low-energy Bose fluid was theoretically studied in different frameworks. In Ref. [23], this problem was studied for an incompressible superfluid, where the spreading of coherence is associated with the decay of a random tangle of quantized vortices (variously known as the Kibble’s vortex tangle [22], superfluid turbulence [55], and Vinen turbulence [26, 33]) and $\ell$ is set by the typical distance between the vortex lines. For $\ell\gg\xi$, where $\xi$ is the size of the vortex core, the prediction is that
$\frac{\textrm{d}\ell}{\textrm{d}t}\propto\frac{\ln{[A\ell/\xi]}}{\ell}\,$, (1)
where $A$ is a dimensionless constant. On the other hand, for coarsening of wave excitations, corresponding to a compressible-fluid flow, approximate kinetic equations give
$\ell(t)\propto t^{\beta}\,$, (2)
with $\beta=1/2$ [21, 39, 41, 34]. In our weakly interacting gas, with the vortex-core size set by the healing length $\xi=1/\sqrt{8\pi na}$, both vortices and waves could play a significant role. Note, however, that the predictions of Eqs. (1) and (2) differ only in a logarithmic correction. Moreover, our measurements of $n_{k}$ probe $\ell$ independently of what type of excitations are dominant and limit its value.
For coarsening fully characterised by the growth of $\ell$ [25], the low-$k$ momentum distribution exhibits self-similar evolution:
$n_{k}(k,t)=\ell^{d}(t)\,f\left[k\ell(t)\right]\,$, (3)
where $d=3$ is the system dimensionality and $f$ is a dimensionless scaling function. In this case $\ell(t)$ is $\propto n_{0}^{1/3}(t)$, where $n_{0}\equiv n_{k}(k=0)$, and the condensate occupation is $N_{0}=(2\pi)^{3}n_{0}/V$. This scaling does not define the absolute value of $\ell$, and here we define it so that in equilibrium $\ell^{3}$ is equal to the system volume; to convert the observed $n_{0}$ values to $\ell$ we also take into account the finite $k$-space resolution of our time-of-flight measurements (Methods). We observe good agreement with Eqs. (2, 3), without the need to invoke the logarithmic correction in Eq. (1), but this does not exclude its relevance on much larger lengthscales; also note that the importance of this correction depends on the unknown value of $A$.
To compare our data to Eq. (2), we first note that it is formally valid only for $t,\ell\rightarrow\infty$. However, one can observe the same scaling for finite $t$ and $\ell$ by absorbing all the effects of the non-universal initial dynamics into time offsets such as seen in Fig. 1b, i.e., by shifting $t\rightarrow t-t^{*}$, where $t^{*}$ depends on the initial state [49] (see also  [56]).
For $\beta=1/2$, in the scaling regime $\ell\propto(t-t^{*})^{1/2}$, and in Fig. 1c we plot $\ell^{2}(t)$. This reveals both the scaling regime where $\ell^{2}$ grows linearly, at a rate that does not depend on the initial state (dashed lines), and the offsets $t^{*}$ for P1,2,3; see also Extended Data Fig. 3. As the system approaches equilibrium, for $\ell^{2}\gtrsim 900\,\upmu{\rm m}^{2}\approx(V/2)^{2/3}$ the scaling breaks down due to finite-size effects; from hereon we focus on the regime before this breakdown.
In Fig. 1d, using the $t^{*}$ values from Fig. 1c, we show that in the scaling regime the low-$k$ momentum distributions for all three initial states can indeed be collapsed onto the same universal curve according to Eq. (3) with $\ell\propto(t-t^{*})^{1/2}$ (see also Extended Data Fig. 4).
We now turn to varying the strength of the interactions that drive the coarsening, and study how this affects the coarsening rate. For $\ell\propto(t-t^{*})^{\beta}$, we define the ‘speed of coarsening’ as $D\equiv\textrm{d}\ell^{1/\beta}/\textrm{d}t$, which is (for any $\beta$) time-invariant in the scaling regime and does not depend on the non-universal $t^{*}$. For a low-energy Bose gas described by the Gross–Pitaevskii equation, the interaction-set units of length and time are $\xi$ and $t_{\xi}\equiv m\xi^{2}/\hbar$, where $\hbar$ is the reduced Planck’s constant and $m$ the atom mass. Hence, on dimensional grounds $\ell/\xi\propto[(t-t^{*})/t_{\xi}]^{\beta}$ and we get
[…]
Specifically for $\beta=1/2$, as observed in Fig. 1, this has a counter-intuitive implication that $D=\textrm{d}\ell^{2}/\textrm{d}t\sim\hbar/m$ does not depend on the interaction strength $na$ (see also [21, 24]). In the hydrodynamic theory of Ref. [23], the same result emerges if one neglects the logarithmic correction in Eq. (1) (which depends on the interactions through $\xi$), because then $\textrm{d}\ell^{2}/\textrm{d}t$ can only be set by the quantum of velocity circulation associated with a quantum vortex, $\kappa=2\pi\hbar/m$.
Figure 3: Universality of the coarsening speed $D$. Our measurements for various interaction strengths show no systematic variation of $D=\textrm{d}\ell^{2}/\textrm{d}t$ (in the scaling regime), and give a combined estimate $D=3.4(3)\,\hbar/m$ (solid line and shading). The four data sets shown in Fig. 1c or Fig. 2b are represented here by the corresponding coloured symbols. The purple triangle indicates a measurement where we reduced the gas density $n$ by a factor of $4.2$, which reduces $T_{\textrm{c}}\propto n^{2/3}$ and the equilibrium condensed fraction (from about 6060% to about 3030%), but $D$ is not affected. The green pentagon indicates a measurement where we instead reduced the system volume $V$ by a factor of $3.5$; in this case $\ell^{2}$ saturates at a correspondingly lower value ($\propto V^{2/3}$), but again $D$ is not affected. For details on the data taken with reduced $n$ or $V$ see Extended Data Fig. 5. The data points corresponding to the same nana values are slightly offset horizontally for visual clarity.
Figure 4: Reaching the coarsening speed limit. a, The growth of $\ell^{2}$ for the same initial state (P1) and gas density, but different $a$. All curves start at $t=0$ and plotting them versus $t-t^{*}$ (with $a$-dependent $t^{*}$) collapses them in the universal scaling regime. The dashed line shows $\ell^{2}=D(t-t^{*})$. For weaker interactions the system approaches this speed limit slower, and reaches it at a larger $\ell$; the solid lines show exponential fits to the early-time data. b, When expressed in the interactions-set units of length, $\xi=1/\sqrt{8\pi na}$, and time, $t_{\xi}\equiv m\xi^{2}/\hbar$, all our data for different Pi, $a$, $V$, and $n$ (see Fig. 3) collapse onto a single curve, meaning that the speed limit is always reached at the same $\ell/\xi$. The solid line shows exponential growth with a time constant $\tau=56\,t_{\xi}$ and the dashed line has slope $mD/\hbar=3.4$. c, Numerically differentiating the data in b, we eliminate the non-universal $t^{*}$ and show as a function of $(\ell/\xi)^{2}$ how $\textrm{d}\ell^{2}/\textrm{d}t$ approaches the universal $D=3.4\,\hbar/m$ and then stops growing; the solid and dashed lines are the same functions as in b. For weaker interactions (larger $\xi$), observing this speed limit requires a larger physical system.
In Fig. 2 we show data for coarsening at $a=100\,a_{0}$ and $400\,a_{0}$, starting from the same initial state P1; the $100\,a_{0}$ data here is the same as in Fig. 1. The images in Fig. 2a show what one intuitively expects – for larger $a$ the condensate emerges sooner. However, in Fig. 2b, plotting $\ell^{2}(t)$ reveals that the effects of the interaction strength are, just like the initial-state effects, confined to the difference in the non-universal $t^{*}$ – in the scaling regime (solid symbols), the rate of the linear growth of $\ell^{2}$ is essentially the same for both $a$.
We performed such measurements for various scattering lengths, and also varied the gas density and system size. In each case we fit the slope $D=\textrm{d}\ell^{2}/\textrm{d}t$ in the scaling regime, and summarise our results in Fig. 3. We observe no systematic variation of $D$ with the interaction strength $na$ and get a combined estimate $D=3.4(3)\,\hbar/m$; here the error in $D$ is purely statistical, while the uncertainty in our box volume leads to a systematic error of ± $\pm\,0.5\,\hbar/m$ (Methods).
Finally, in Fig. 4 we study how the system approaches the long-time coarsening speed $D$, which also reconciles our observations with the finite-system intuition (and experience) that for stronger interactions condensates form faster, and for $a\rightarrow 0$ they do not form at all.
As we illustrate in Fig. 4a for fixed $n$, the same initial state (P1), and three values of $a$, for weaker interactions the gas takes longer to join the universal scaling trajectory, and joins it at a larger value of $\ell$. Here all curves start at $t=0$ and plotting versus $t-t^{*}$ collapses them in the scaling regime; the dashed line shows $\ell^{2}=D(t-t^{*})$.
In Fig. 4b we show that expressing all our data for different Pi, $a$, $V$, and $n$ (see Fig. 3) in terms of $\xi$ and $t_{\xi}$ collapses them onto a single curve. Here, the dashed line has slope $mD/\hbar=3.4$ and the solid line that captures the approach to the scaling regime is an exponential with a time constant $\tau=56\,t_{\xi}$ (see also Extended Data Fig. 6). Numerically differentiating these data, and thus eliminating the non-universal $t^{*}$, in Fig. 4c we show how $\textrm{d}\ell^{2}/\textrm{d}t$ approaches the universal $D=3.4\,\hbar/m$ (dashed line) as a function of $(\ell/\xi)^{2}$.
The dimensionless results in Figs. 4b and 4c imply that for any interaction strength the system would eventually, for $\ell\gg\xi$, exhibit the same coarsening speed $D$. However, the system size required to observe this is larger for larger $\xi$ (smaller $na$), and diverges for $na\rightarrow 0$. As we discuss in Methods and Extended Data Fig. 7, previous experiments on the emergence of extended coherence during far-from-equilibrium condensation, in both harmonic [14] and box [46] traps, were not in the universal-speed regime; consequently, for tunable interactions in Ref. [46], the observed relaxation time was $\propto 1/a$.
Our results should be relevant across many fields, from cold atoms and the conventional low-temperature physics [26, 33] to cosmology [22, 27] and high-energy physics [40]. The fact that $\hbar$ and $m$ appear in $D$ only through their ratio, or the quantum of circulation $\kappa$, implies that the results are also applicable to systems where the underlying physics is not quantum [57]. They should also be relevant for benchmarking the theories of ultrarelativistic systems, for which $\beta=1/2$ is also predicted [39], but in that case the effective mass, $m_{\textrm{eff}}\sim\hbar/(\xi c)$ (where $c$ is the speed of light), and hence $D\sim\xi c$, depend on the interactions.
The value of $D=3.4\,\hbar/m$, equal to $5.5\,\upmu{\rm m}^{2}/{\rm ms}$ for 39K, has curious implications for the emergence of coherence on truly macroscopic lengthscales. For example, for coherence to spread via coarsening over 11\,cm would require hours, and even possible logarithmic corrections cannot change this conclusion significantly. However, an interesting question is whether this speed limit can be broken by a fundamentally different preparation protocol, for example by melting a Mott insulator (through a quantum phase transition) to obtain a superfluid with long-range coherence.
In the future, it would be interesting to disentangle the roles of waves and vortices during coarsening, by directly imaging the latter, to search for possible logarithmic corrections to the coarsening speed, and to perform similar measurements for fermionic superfluids and gases with long-range interactions.
We thank Andrey Karailiev, Maciej Gałka, Timon Hilker, Mikhail Lukin, Robert Smith, Tobias Donner, Gregory Eyink, and Vladimir Rosenhaus for useful discussions and comments on the manuscript. This work was supported by EPSRC [Grants No. EP/P009565/1 and EP/Y01510X/1], ERC (UniFlat), and STFC [Grants No. ST/T006056/1 and No. ST/Y004469/1]. C. E. acknowledges support from Jesus College (Cambridge). Z. H. acknowledges support from the Royal Society Wolfson Fellowship.
Author contributions  G.M. led the project and collected the data. G.M., M.G., J.E., C.E., and Z.H. conceptualised the project. All authors contributed to the experimental setup, data analysis, and writing of the manuscript.
Competing interests The authors declare no competing interests.